# Out of Distribution Detection for Efficient Continual Learning in Quality Prediction for Arc Welding

### v1

@@ Abstract
Modern manufacturing relies heavily on fusion welding processes, including gas metal arc welding (GMAW). Despite significant advances in machine learning-based quality prediction, current models exhibit critical limitations when confronted with the inherent distribution shifts that occur in dynamic manufacturing environments. In this work, we extend the VQ-VAE Transformer architecture—previously demonstrating state-of-the-art performance in weld quality prediction—by leveraging its autoregressive loss as a reliable out-of-distribution (OOD) detection mechanism. Our approach exhibits superior performance compared to conventional reconstruction methods, embedding error-based techniques, and other established baselines. By integrating OOD detection with continual learning strategies, we optimize model adaptation, triggering updates only when necessary and thereby minimizing costly labeling requirements. We introduce a novel quantitative metric that simultaneously evaluates OOD detection capability while interpreting in-distribution performance. Experimental validation in real-world welding scenarios demonstrates that our framework effectively maintains robust quality prediction capabilities across significant distribution shifts, addressing critical challenges in dynamic manufacturing environments where process parameters frequently change. This research makes a substantial contribution to applied artificial intelligence by providing an explainable and at the same time adaptive solution for quality assurance in dynamic manufacturing processes - a crucial step towards robust, practical AI systems in the industrial environment.
AL
arc length
VQ-VAE
vector quantised variational autoencoder
VQ
vector quantisation
VAE
variational autoencoder
CNN
convolutional neural network
DIN
German Institute for Standardisation Registered Association
DNP
distance from nozzle to plate
EWC
elastic weight consolidation
GFR
gas flow rate
GMAW
gas metal arc welding
HIVE-COTE
LSTM
long short-term memory
MAG
metal active gas
MAS
memory-aware synapses
MIG
metal inert gas
RNN
recurent neural network
VPRS
variable precision rough set
WFR
wire feed rate
WS
welding speed
ML
machine learning
DL
deep learning
MLP
multilayer perceptron
GRU
gated recurrent unit
CV
computer vision
NLP
natural language processing
CNN
convolutional neural networks
CMT
cold metal transfer
VQ-VAE Tr
VQ-VAE Transformer
@@ 1 Introduction
Arc welding processes, particularly GMAW, form the backbone of modern manufacturing across automotive, marine, and aerospace industries [2]. Despite its widespread adoption, GMAW presents significant challenges stemming from the complex interplay between process parameters—current, voltage, welding speed—and resultant weld quality. Traditional quality assessment approaches rely heavily on operator expertise and post-process inspection, limiting real-time quality prediction capabilities. These challenges can be effectively addressed through machine learning and deep learning methodologies, which leverage sensor and process data to develop predictive models capable of real-time quality assessment.
@@
This work addresses the critical challenge of detecting out-of-distribution (OOD) data relative to a model’s learned distribution, particularly during transitions between fundamentally different welding types (e.g., overlap to T-joints). Our work builds upon an existing state-of-the-art VQ-VAE Transformer (VQ-VAE Tr) model for arc welding quality prediction [11]. While this model has advantages in the manufacturing domain through a combination of unsupervised and supervised learning, the architecture offers a high potential for OOD detection. We present an approach that leverages autoregressive loss as an effective OOD indicator, outperforming traditional methods (e.g. Maximum Softmax Probability (MSP) [14] or ODIN [19]) as well as reconstruction- and quantization error-based detection approaches. This OOD detection mechanism serves as a trigger for initiating continual learning only when necessary, balancing adaptation needs with computational efficiency, reducing the costs of labeling data and using training resources.
@@
    We demonstrate that autoregressive loss of a VQ-VAE Tr architecture serves as a superior OOD indicator compared to reconstruction or quantization errors and other baselines, providing a reliable trigger mechanism for model adaptation without requiring additional computational overhead.
@@ 2.1 Quality Prediction in Welding
Welding quality prediction has evolved from traditional mathematical models requiring manual feature engineering to modern deep learning models that automatically learn representations from sensor data [42, 11]. Recent advances in machine learning, including neural networks of varying complexity and recurrent architectures, have improved predictive accuracy for welding performance [40, 17, 43, 41]. Among these, the VQ-VAE Tr architecture introduced by [11] achieved state-of-the-art performance in time series welding quality prediction by combining unsupervised and supervised learning approaches.
@@
Our work addresses this gap by extending the VQ-VAE Tr architecture, leveraging its inherent but previously untapped potential for distribution shift detection through its multi-faceted representational capabilities. We develop robust OOD detection mechanisms that enable the system to identify manufacturing process variations and trigger appropriate model adaptation when necessary.
@@ 2.2.3 Parametrized approaches
Our work builds upon the generative modeling paradigm but addresses these limitations by leveraging the autoregressive loss of the transformer component within our VQ-VAE Tr architecture. Similar to approaches in medical imaging [9], we exploit the inherent temporal modeling capabilities as a natural OOD indicator. This eliminates the need for separate detection modules while providing superior sensitivity to the temporal dependencies characteristic of welding process data, effectively bridging the gap between unsupervised OOD detection and supervised quality prediction tasks.
@@ 3.1 Gas Metal Arc Welding
Arc welding is a specific type of fusion welding used to join metallic objects [35]. In contrast to other welding techniques like gas or laser welding, arc welding utilizes electrical energy to create the heat necessary to melt welding wire and workpiece. Our use case lies in the field of GMAW, a welding technique that uses a continuously fed welding wire as the electrode and shielding gas to protect the material from atmospheric influences.
@@ 4.3 ROC-based Threshold Optimization
Our implementation proceeds through four discrete stages. First, we partition the validation set into correctly and incorrectly classified instances by comparing predicted labels against ground truth annotations. Second, we compute appropriate OOD indicators for each instance (e.g., MSP). For discriminative models, we extract confidence scores derived from maximum softmax probabilities, whereas for generative architectures like our VQ-VAE Tr, we utilize error metrics—specifically autoregressive loss.
@@ 5.1 Out-of-Distribution Detection
We evaluated the proposed OOD detection approach against several established baselines across both welding quality prediction and OOD detection capabilities. The baseline models included: (i) CODiT, a temporal-based OOD detection framework leveraging conformity scores; (ii) a standard multi-layer perceptron (MLP); (iii) MLP with evidential deep learning (EDL); and (iv) a VQ-VAE-based MLP classifier. For OOD detection methods, we compared Maximum Softmax Probability (MSP), ODIN, Vacuity-based detection, and Mahalanobis distance approaches. Our primary model, the VQ-VAE Tr, was implemented following the architecture and training routine described in [11].
@@ 5.2 OOD Results
The results presented in Table 2 demonstrate that the VQ-VAE Tr architecture achieves superior performance on the primary welding quality prediction task, achieving an F1-score that is 16% higher than the second-best model. This represents the best classification performance of all the models evaluated.
The critical finding of our evaluation concerns the OOD detection capabilities across different approaches. The VQ-VAE Tr using autoregressive loss as the detection signal achieves the highest OOD detection F1-score of 0.35 ± 0.08, followed closely by the Vacuity-based approach at 0.34 ± 0.02. Notably, traditional reconstruction and encoder-based OOD detection methods within the VQ-VAE framework yielded substantially lower performance ($\mathcal{S}_{\text{recon}}$=0.28 ± 0.07 and $\mathcal{S}_{\text{quant}}$=0.20 ± 0.08, respectively). This phenomenon can be attributed to the VQ-VAE’s generalization capabilities across distribution shifts, which manifest in robust reconstruction performance but fail to transfer to downstream classification tasks. While the encoder-decoder architecture demonstrates resilience to distributional variations, the classification head remains susceptible to these shifts, highlighting the differential impact of OOD data on distinct model components.
@@ 6 Conclusion and Outlook
In this work, we addressed a fundamental challenge in industrial machine learning: maintaining robust predictive performance while minimizing costly labeling requirements as manufacturing processes evolve. We present a novel approach to out-of-distribution detection in dynamic welding processes, demonstrating that autoregressive loss from a VQ-VAE Tr architecture provides a reliable mechanism for identifying distribution shifts in welding quality prediction tasks. By integrating OOD detection with selective continual learning.
@@
By enabling autonomous detection and adaptation to process changes, this framework paves the way for self-maintaining quality systems that remain robust across welding configurations, parameter drifts, and unexpected distribution shifts. Future research directions include extending the approach to multi-modal sensor data, investigating more sophisticated continual learning strategies that leverage the discrete token representations, and exploring the application of this framework to other manufacturing processes beyond welding. While demonstrated on the VQ-VAE Tr, the core principle—leveraging autoregressive loss as an OOD indicator to trigger selective adaptation—could potentially enhance other autoregressive models and manufacturing domains where temporal patterns encode critical process dynamics.

### v2

@@ Abstract
Modern manufacturing relies heavily on fusion welding processes, including gas metal arc welding. Despite significant advances in machine learning-based quality prediction, current models exhibit critical limitations when confronted with the inherent distribution shifts that occur in dynamic manufacturing environments. In this work, we extend the VQ-VAE Transformer architecture—previously demonstrating state-of-the-art performance in weld quality prediction—by leveraging its autoregressive loss as a reliable out-of-distribution (OOD) detection mechanism. Our approach exhibits superior performance compared to conventional reconstruction methods, embedding error-based techniques, and other established baselines. By integrating OOD detection with continual learning strategies, we optimize model adaptation, triggering updates only when necessary and thereby minimizing costly labeling requirements. We introduce a novel quantitative metric that simultaneously evaluates OOD detection capability while interpreting in-distribution performance. Experimental validation in real-world welding scenarios demonstrates that our framework effectively maintains robust quality prediction capabilities across significant distribution shifts, addressing critical challenges in dynamic manufacturing environments where process parameters frequently change. This research makes a substantial contribution to applied artificial intelligence by providing an explainable and at the same time adaptive solution for quality assurance in dynamic manufacturing processes - a crucial step towards robust, practical AI systems in the industrial environment.
@@ 1 Introduction
Arc welding processes, particularly gas metal arc welding, form the backbone of modern manufacturing across automotive, marine, and aerospace industries [2]. Despite its widespread adoption, gas metal arc welding presents significant challenges stemming from the complex interplay between process parameters—current, voltage, welding speed—and resultant weld quality. Traditional quality assessment approaches rely heavily on operator expertise and post-process inspection, limiting real-time quality prediction capabilities. These challenges can be effectively addressed through machine learning and deep learning methodologies, which leverage sensor and process data to develop predictive models capable of real-time quality assessment.
@@
This work addresses the critical challenge of detecting out-of-distribution (OOD) data relative to a model’s learned distribution, particularly during transitions between fundamentally different welding types (e.g., overlap to T-joints). Our work builds upon an existing state-of-the-art VQ-VAE Transformer model for arc welding quality prediction [12]. While this model has advantages in the manufacturing domain through a combination of unsupervised and supervised learning, the architecture offers a high potential for OOD detection. We present an approach that leverages autoregressive loss as an effective OOD indicator, outperforming traditional methods (e.g. Maximum Softmax Probability (MSP) [14] or ODIN [19]) as well as reconstruction- and quantization error-based detection approaches. This OOD detection mechanism serves as a trigger for initiating continual learning only when necessary, balancing adaptation needs with computational efficiency, reducing the costs of labeling data and using training resources.
@@
    We demonstrate that autoregressive loss of a VQ-VAE Transformer architecture serves as a superior OOD indicator compared to reconstruction or quantization errors and other baselines, providing a reliable trigger mechanism for model adaptation without requiring additional computational overhead.
@@ 2.1 Quality Prediction in Welding
Welding quality prediction has evolved from traditional mathematical models requiring manual feature engineering to modern deep learning models that automatically learn representations from sensor data [42, 12]. Recent advances in machine learning, including neural networks of varying complexity and recurrent architectures, have improved predictive accuracy for welding performance [40, 17, 43, 41]. Among these, the VQ-VAE Transformer architecture introduced by [12] achieved state-of-the-art performance in time series welding quality prediction by combining unsupervised and supervised learning approaches.
@@
Our work addresses this gap by extending the VQ-VAE Transformer architecture, leveraging its inherent but previously untapped potential for distribution shift detection through its multi-faceted representational capabilities. We develop robust OOD detection mechanisms that enable the system to identify manufacturing process variations and trigger appropriate model adaptation when necessary.
@@ 2.2.3 Parametrized approaches
Our work builds upon the generative modeling paradigm but addresses these limitations by leveraging the autoregressive loss of the transformer component within our VQ-VAE Transformer architecture. Similar to approaches in medical imaging [9], we exploit the inherent temporal modeling capabilities as a natural OOD indicator. This eliminates the need for separate detection modules while providing superior sensitivity to the temporal dependencies characteristic of welding process data, effectively bridging the gap between unsupervised OOD detection and supervised quality prediction tasks.
@@ 3.1 Gas Metal Arc Welding
Arc welding is a specific type of fusion welding used to join metallic objects [35]. In contrast to other welding techniques like gas or laser welding, arc welding utilizes electrical energy to create the heat necessary to melt welding wire and workpiece. Our use case lies in the field of gas metal arc welding, a welding technique that uses a continuously fed welding wire as the electrode and shielding gas to protect the material from atmospheric influences.
@@ 4.3 ROC-based Threshold Optimization
Our implementation proceeds through four discrete stages. First, we partition the validation set into correctly and incorrectly classified instances by comparing predicted labels against ground truth annotations. Second, we compute appropriate OOD indicators for each instance (e.g., MSP). For discriminative models, we extract confidence scores derived from maximum softmax probabilities, whereas for generative architectures like our VQ-VAE Transformer, we utilize error metrics—specifically autoregressive loss.
@@ 5.1 Out-of-Distribution Detection
We evaluated the proposed OOD detection approach against several established baselines across both welding quality prediction and OOD detection capabilities. The baseline models included: (i) CODiT, a temporal-based OOD detection framework leveraging conformity scores; (ii) a standard multi-layer perceptron (MLP); (iii) MLP with evidential deep learning (EDL); and (iv) a VQ-VAE-based MLP classifier. For OOD detection methods, we compared Maximum Softmax Probability (MSP), ODIN, Vacuity-based detection, and Mahalanobis distance approaches. Our primary model, the VQ-VAE Transformer, was implemented following the architecture and training routine described in [12].
@@ 5.2 OOD Results
The results presented in Table 2 demonstrate that the VQ-VAE Transformer architecture achieves superior performance on the primary welding quality prediction task, achieving an F1-score that is 16% higher than the second-best model. This represents the best classification performance of all the models evaluated.
The critical finding of our evaluation concerns the OOD detection capabilities across different approaches. The VQ-VAE Transformer using autoregressive loss as the detection signal achieves the highest OOD detection F1-score of 0.35 ± 0.08, followed closely by the Vacuity-based approach at 0.34 ± 0.02. Notably, traditional reconstruction and encoder-based OOD detection methods within the VQ-VAE framework yielded substantially lower performance ($\mathcal{S}_{\text{recon}}$=0.28 ± 0.07 and $\mathcal{S}_{\text{quant}}$=0.20 ± 0.08, respectively). This phenomenon can be attributed to the VQ-VAE’s generalization capabilities across distribution shifts, which manifest in robust reconstruction performance but fail to transfer to downstream classification tasks. While the encoder-decoder architecture demonstrates resilience to distributional variations, the classification head remains susceptible to these shifts, highlighting the differential impact of OOD data on distinct model components.
@@ 6 Conclusion and Outlook
In this work, we addressed a fundamental challenge in industrial machine learning: maintaining robust predictive performance while minimizing costly labeling requirements as manufacturing processes evolve. We present a novel approach to out-of-distribution detection in dynamic welding processes, demonstrating that autoregressive loss from a VQ-VAE Transformer architecture provides a reliable mechanism for identifying distribution shifts in welding quality prediction tasks. By integrating OOD detection with selective continual learning.
@@
By enabling autonomous detection and adaptation to process changes, this framework paves the way for self-maintaining quality systems that remain robust across welding configurations, parameter drifts, and unexpected distribution shifts. Future research directions include extending the approach to multi-modal sensor data, investigating more sophisticated continual learning strategies that leverage the discrete token representations, and exploring the application of this framework to other manufacturing processes beyond welding. While demonstrated on the VQ-VAE Transformer, the core principle—leveraging autoregressive loss as an OOD indicator to trigger selective adaptation—could potentially enhance other autoregressive models and manufacturing domains where temporal patterns encode critical process dynamics.